# Don’t be Fooled: The Misinformation Effect of Explanations in Human-AI Collaboration

### v1

@@ 5.2. RQ2: Impact of Incorrect Explanations on Reasoning
To analyze participants’ reasoning capabilities as they transition from knowing ”how” to understanding ”why”, we assess the answers they provided in the open-text questionnaire. The primary focus is on evaluating how the accuracy of AI-generated explanations influences the participants’ reasoning abilities to understand how incorrect explanations impact their knowledge.
We assess participants’ reasoning abilities following the procedure of Huang et al. (2024). By doing so, we use LLMs to evaluate the correctness of participants’ answers and compute a reasoning score: zero represents a completely incorrect or irrelevant answer, and five represents a completely correct and comprehensive answer. The scores are then analyzed using independent t-tests to compare the performance between treatments, and the Holm-Bonferroni method is used to correct them.
In the incorrect explanation treatment (), participants score lower on average than those in the control group (). However, the difference is not significant (). When comparing the incorrect explanation treatment to the AI treatment(), participants exposed to incorrect explanations perform significantly worse (). The comparison between the incorrect explanation treatment and the correct explanation treatment () shows similar results (). This suggests that incorrect explanations tend to decrease reasoning abilities compared to correct explanations or only AI advice. These results highlight the critical impact of explanation accuracy on reasoning, with incorrect explanations leading to notable decreases in reasoning performance, particularly when compared to AI-generated or correct explanations.
@@ 5.3. RQ3: Impact of Incorrect Explanations on Human-AI Team Performance
Analog to the post-test performance, we run a regression analysis to reveal the effect of confounding factors on the main task performance (see Table 7 in Appendix A). The model investigating the main task performance explains of the variance (), showing that participants in AI-supported treatments outperform those in the control group. Participants in the correct explanation treatment demonstrate the most substantial improvement over the control group, with a significant positive effect (). This is closely followed by the AI treatment () and the incorrect explanation treatment (). Furthermore, AI trust shows a significant negative association with performance (), indicating a possible nuanced effect of trust on how participants interact with the AI.
These findings suggest that while incorrect explanations can aid immediate task performance, they may hinder the retention of procedural knowledge (as the post-test performance is below the main-task performance). We use a repeated measures approach and run a mixed-effects model to analyze how procedural knowledge is affected by first providing and then removing AI support. We define performance as the dependent variable, the type of AI support as the independent variable, the different stages—with AI support in the main task and without AI support in the post-test—as the mediating factor, cognitive load, AI trust, AI knowledge and AI usefulness as control factors and participant ID as a random factor.
@@ 6.1. Summary of Findings
With the rise of AI in decision-making domains, it is crucial to understand how the interaction with AI affects decision-makers. Prior research so far has either focused on the implications of incorrect AI advice (e.g., (Schemmer et al., 2023c; Bansal et al., 2021)) or explored how the correctness of explanations affects trust and reliance on AI (e.g., (Morrison et al., 2024; Cabitza et al., 2024)). This work investigates how the correctness of AI explanations impacts humans through collaboration with an AI. In a classification task, participants were assigned to one of four treatments: no AI support, AI advice only, AI advice with correct explanations, and AI advice with incorrect explanations. We measured the task performance before, during, and post-collaboration to derive humans’ procedural knowledge, reasoning, and the human-AI team performance. The results show that all AI-assisted treatments led to significantly higher human-AI team performance during the main task compared to the control group, demonstrating the merits of AI support. However, the correctness of the explanations played a crucial role: Participants who had received correct explanations exhibited the highest main task performance, followed by those who received only AI advice without explanations, dominating those who had received incorrect explanations. Importantly, the findings reveal that the accuracy of the explanations has a lasting impact on procedural knowledge development. While incorrect explanations temporarily increase performance during the main task, the procedural knowledge was impaired post-collaboration, ultimately hindering knowledge retention compared to correct explanations or no explanations. Moreover, these results must be considered with caution. Even though there was an increase in procedural knowledge, participants’ reasoning capabilities in the incorrect explanation treatment were below the ones from the control group. This indicates that incorrect explanations can undermine the durable benefits of human-AI collaboration despite initial performance improvements. These results contribute to the HCI literature by underscoring the nuanced effects of explanation accuracy on human-AI collaboration. The findings emphasize the importance of designing AI-based support that provides accurate explanations to not only enhance human decision-making but also maintain and improve humans’ procedural knowledge and reasoning in the long run. Overall, the study’s key takeaway is that, while AI support can generally improve performance, the correctness of the explanations provided by the AI is a crucial determinant of human-AI collaboration, influencing humans’ ability to draw conclusions and perform the task autonomously post-collaboration with an AI. These insights can inform the design of more effective and human-centric AI-based decision support systems.
@@ 6.2. Implications
This work makes several contributions to the field of HCI by deepening the understanding of how the correctness of AI explanations influences humans’ procedural knowledge and reasoning ability, as well as the human-AI team performance. While previous research has often focused on the benefits of AI assistance in enhancing decision-making (Schemmer et al., 2023a), this study offers a more nuanced perspective, highlighting the critical role of explanation correctness on humans’ knowledge and understanding. It complements the large corpus of HCI literature on XAI (i.e.; (Morrison et al., 2024; Schoeffer et al., 2024; Schemmer et al., 2023c; Hemmer et al., 2023; Spitzer et al., 2024b)) by explicitly investigating the influence of incorrect explanations for scenarios when the AI advice is correct and identifies the negative repercussions for humans.
@@
Taking a human-centric perspective. In our study, we could see that the human-AI team performance improved over the participants’ initial task performance. This pattern, also observed by prior research (Inkpen et al., 2023; Hemmer et al., 2024), showcases the potential of humans collaborating with AI. Even though the scenario in which the human-AI team performance exceeds the performance of human or AI alone—also referred to as complementary team performance (Bansal et al., 2019)—could not be reached (due to the study design choices this was not feasible as the AI advice was always correct), our findings showcase the two sides of explanations: while correct explanations improve the human-AI team performance compared to only receiving AI advice (bright side) and demonstrate the merits of XAI, incorrect explanations decrease the human-AI team performance compared to only receiving AI advice (dark side) and outline the pitfalls of XAI. Thus, it is important to implement mechanisms that allow humans to verify the correctness of explanations, for instance, through reflection mechanisms (Ehsan and Riedl, 2024). Furthermore, we could also identify AI trust as a confounding factor during the main task and cognitive load during the post-test. While these findings align with prior research (Ueno et al., 2022; Buçinca et al., 2021; Westphal et al., 2023), they also emphasize the important role of individuals’ characteristics and traits in human-AI collaboration. Therefore, research has to anticipate these factors in the design of robust and safe explanations.
@@
Anchoring on explanations. We also take a critical view on the phenomenon prior research has identified in collaboration scenarios between humans and AI in which the AI provides explanations: the anchoring effect on explanations (Wang et al., 2019; Bansal et al., 2021). The effect occurs when human decision-makers fixate on the explanation and form an incorrect understanding. In our study, participants in the incorrect explanation treatment demonstrated this behavior by achieving lower procedural knowledge on the post-test compared to the other AI-supported treatments and also by showing lower reasoning capabilities than participants in the control group. As a result, they made the right decisions for the wrong reasons, a condition often referred to as the Clever Hans Effect. (Schramowski et al., 2020; Kauffmann et al., 2020). Although research has adopted the term to describe AI behavior, in our study, humans demonstrate the effect. While this illustrates a major incision in the human knowledge structure, it is crucial for research and practice to develop mechanisms to counteract this effect.
@@
Design guidelines for practice. Our findings confirm that AI support enhances immediate task performance, even when the explanations are incorrect. Even though this observation aligns with prior HCI studies demonstrating that explanations can increase human-AI team performance, such as those by Bansal et al. (2019) and Schemmer et al. (2023c), we extend prior research by revealing the negative impacts of explanations. Our work shows that the performance boost provided by AI during the collaboration can be short-term if the explanations are incorrect, as they negatively impact humans’ procedural knowledge and reasoning post-collaboration. This finding advances the existing literature on XAI, which has primarily emphasized factors like transparency, reliance, and user understanding during the collaboration(Lai and Tan, 2019; Ehsan et al., 2018; Morrison et al., 2024). Our research indicates that while correct explanations foster procedural knowledge development, incorrect explanations can lead to the erosion of knowledge over time, even when they temporarily enhance task performance in the collaboration with AI. The findings also underscore the importance of designing AI that prioritize the correctness of explanations, especially in contexts where AI support is not always provided and human decision-makers have to be able to decide autonomously, for instance, as required by the EU AI act (Commission, 2021). We suggest that AI should integrate mechanisms allowing humans to evaluate the correctness of explanations critically. For instance, incorporating counterfactual explanations, as explored by previous research (Wachter et al., 2017; Goyal et al., 2019; Schemmer et al., 2023b), could help users better understand the AI’s decision-making process and mitigate the adverse effects of incorrect explanations. Similarly, maintaining the human ability to perform tasks autonomously without AI support is favorable at the workplace. The role of XAI can impact the transfer of knowledge to counter risks of demographic change (Spitzer et al., 2024b). Incorrect explanations impairing knowledge development have the potential to endanger the successful business operations of organizations. Therefore, we advise designers to incorporate not only mechanisms to detect incorrect explanations post-hoc (Xu et al., 2023; Zhou et al., 2023; Hartwig et al., 2023) but also in the design of the AI.
@@
The impact for organizations. Prior research has mainly explored effects that occur during the collaboration of humans and AI. For instance, Schemmer et al. (2023c) conceptualize the relationship between the appropriateness of reliance and how it relates to the human-AI team performance. Morrison et al. (2024) advance this view by the dimension of explanations and explore the effects of their correctness on humans’ appropriate AI reliance. By outlining potential downsides of XAI, this work addresses the impact of explanations’ correctness on humans post-collaboration. Our study demonstrates that their procedural knowledge and reasoning are impaired when they are provided with incorrect explanations and the AI support is removed. Taking in a human-centric perspective (Horvatić and Lipic, 2021), this repercussion presents harm to not only humans’ individual knowledge development (Bhatt, 2000) but also to their ability to provide meaningful assets to organizations (Nonaka and Takeuchi, 2007; Davenport, 1998). Maintaining individuals as valuable assets to organizations is crucial because it directly influences organizational innovation, efficiency, and adaptability in a competitive market. Sustaining humans’ knowledge development can foster a more resilient and informed workforce capable of driving sustained success (Grant, 1996).
@@ 6.3. Limitations and Future Work
Despite the valuable insights gained from this study, several limitations must be acknowledged, offering avenues for future research. First, the study explores how incorrect explanations affect humans in AI-assisted decision-making. To investigate the impact of such explanations, we designed a study with four treatments of different AI support: no support, only AI advice, AI advice with correct explanations, and AI advice with incorrect explanations. In real-world applications, the interaction with AI that provides only correct or incorrect explanations is rather unlikely. It presents valuable means to take the first steps to investigate the impact of incorrect explanations but does not reflect the real world. Future research could take on this aspect to extend our findings and evaluate how a mix of correct and incorrect explanations—a mix that is more realistic for deployed AI—affects humans’ procedural knowledge and reasoning. Especially, different ratios of the correctness of explanations could provide further insights and advance the field.
@@
Second, the focus of the study was mainly on measuring the task performance to derive insights into humans’ procedural knowledge and inform about the impact on human-AI team performance. However, in real-world scenarios, performance might not be the only metric relevant. Other measures, like appropriate reliance (Schemmer et al., 2023c), trust or fairness (Schoeffer et al., 2024) in the AI, might also be of high relevance in AI-assisted decision-making as previous studies show (Hemmer et al., 2023; Bansal et al., 2021). It is of high importance to explore how these factors change over time and under the effect of incorrect explanations. Exploring the temporal implications can extend the views and offer new insights that support the robust and effective design of AI.
@@
Lastly, the study primarily relied on short-term measures based on task performance, which may not fully capture the long-term impact of AI support on human knowledge and how the correctness of explanations impacts the human-AI team performance. Future research could employ longitudinal designs to assess how incorrect explanations influence procedural knowledge development and reasoning over extended periods and across multiple tasks. This approach would offer a deeper understanding of how different types of explanations contribute to sustained knowledge development, aligning with the principles of human-centered AI.

### v2

@@ 5.2. RQ2: Impact of Incorrect Explanations on Reasoning
To analyze participants’ reasoning capabilities as they transition from knowing ”how” to understanding ”why”, we assess the answers they provided in the open-text questionnaire. The primary focus is on evaluating how the accuracy of AI-generated explanations influences the participants’ reasoning abilities to understand how incorrect explanations impact their knowledge. We provide some randomly selected answers to these open-ended questions in the Appendix A in Table 11. All participants’ results are provided in the supplementary materials.
The scores are then analyzed using independent t-tests to compare the performance between treatments, and the Holm-Bonferroni method is used to correct them.
In the incorrect explanation treatment (), participants score lower on average than those in the control group (). However, the difference is not significant (). When comparing the incorrect explanation treatment to the AI treatment(), participants exposed to incorrect explanations perform significantly worse (). The comparison between the incorrect explanation treatment and the correct explanation treatment () shows similar results (). While not significant at alpha = 0.1, we see a tendency for incorrect explanations to decrease reasoning abilities compared to correct explanations or only AI advice. Participants exposed to incorrect explanations decrease their reasoning ability. With this decrease in the incorrect explanation treatment compared to the correct explanations and AI advice only treatment, there is no significant difference compared to the control group. This means that receiving incorrect explanations has no difference in its impact on participants’ understanding than no AI support. This could mean that the incorrect explanations might still be helpful in certain scenarios to participants to identify actual correct characteristics of the architectural styles.
@@
Additionally, we analyze further control variables that might influence participants’ reasoning capabilities. To do so, we run a regression analysis, similar to Section 5.1, using cognitive load, trust, AI knowledge, and AI usefulness as control variables. We find that AI usefulness has a direct effect on the reasoning capabilities (, ). We report the results in Appendix A in Table 12. Thus, when participants perceive the AI as useful, they also build a better understanding of the underlying task domain. This is an interesting finding that we further discuss in Section 6.
@@ 5.3. RQ3: Impact of Incorrect Explanations on Human-AI Team Performance
Analog to the post-test performance, we run a regression analysis to reveal the effect of confounding factors on the main task performance (see Table 7 in Appendix A). The model investigating the main task performance explains of the variance (), showing that participants in AI-supported treatments outperform those in the control group. Participants in the correct explanation treatment demonstrate the most substantial improvement over the control group, with a significant positive effect (). This is closely followed by the AI treatment () and the incorrect explanation treatment (). Furthermore, AI trust shows a significant negative association with performance (), indicating a possible nuanced effect of trust on how participants interact with the AI. We take further steps and analyze AI trust between the different treatments. AI trust is highest for the correct explanation treatment (), followed by the incorrect explanation treatment (), the control group (), and the AI treatment (). This finding shows that explanations increase trust independently of their correctness. Interestingly, as the score is higher in the control group than in the AI treatment, only providing AI advice does not seem to benefit participants’ trust in the AI.
The findings suggest that while incorrect explanations can aid immediate task performance, they may hinder the retention of procedural knowledge (as the post-test performance is below the main-task performance). We use a repeated measures approach and run a mixed-effects model to analyze how procedural knowledge is affected by first providing and then removing AI support. We define performance as the dependent variable, the type of AI support as the independent variable, the different stages—with AI support in the main task and without AI support in the post-test—as the mediating factor, cognitive load, AI trust, AI knowledge and AI usefulness as control factors and participant ID as a random factor.
@@ 6.1. Summary of Findings
With the rise of AI in decision-making domains, it is crucial to understand how the interaction with AI affects decision-makers. Prior research so far has either focused on the implications of incorrect AI advice (e.g., (Schemmer et al., 2023b; Bansal et al., 2021)) or explored how the correctness of explanations affects trust and reliance on AI (e.g., (Morrison et al., 2024; Cabitza et al., 2024a)). This work investigates how the correctness of AI explanations impacts humans through collaboration with an AI. In a classification task, participants were assigned to one of four treatments: no AI support, AI advice only, AI advice with correct explanations, and AI advice with incorrect explanations. We measured the task performance before, during, and post-collaboration to derive humans’ procedural knowledge, reasoning, and the human-AI team performance. The results show that all AI-assisted treatments led to significantly higher human-AI team performance during the main task compared to the control group, demonstrating the merits of AI support. However, the correctness of the explanations played a crucial role: Participants who had received correct explanations exhibited the highest main task performance, followed by those who received only AI advice without explanations, dominating those who had received incorrect explanations. Importantly, the findings reveal that the accuracy of the explanations has a lasting impact on procedural knowledge development. While incorrect explanations temporarily increase performance during the main task, the procedural knowledge was impaired post-collaboration, ultimately hindering knowledge retention compared to correct explanations or no explanations. Moreover, these results must be considered with caution. Even though there was an increase in procedural knowledge, participants’ reasoning capabilities in the incorrect explanation treatment were below the ones from the control group. This indicates that incorrect explanations can undermine the durable benefits of human-AI collaboration despite initial performance improvements. These results contribute to the HCI literature by underscoring the nuanced effects of explanation accuracy on human-AI collaboration. The findings emphasize the importance of designing AI-based support that provides accurate explanations to not only enhance human decision-making but also maintain and improve humans’ procedural knowledge and reasoning in the long run. Overall, the study’s key takeaway is that, while AI support can generally improve performance, the correctness of the explanations provided by the AI is a crucial determinant of human-AI collaboration, influencing humans’ ability to draw conclusions and perform the task autonomously post-collaboration with an AI. These insights can inform the design of more effective and human-centric AI-based decision support systems.
@@ 6.2. Implications
This work makes several contributions to the field of HCI by deepening the understanding of how the correctness of AI explanations influences humans’ procedural knowledge and reasoning ability, as well as the human-AI team performance. Although previous research has often focused on the benefits of AI assistance in enhancing decision-making (Schemmer et al., 2023a), this study offers a more nuanced perspective, highlighting the critical role of explanation correctness on humans’ knowledge and understanding. It complements the large corpus of HCI literature on XAI (i.e., (Morrison et al., 2024; Schoeffer et al., 2024; Speith, 2022)) by explicitly investigating the influence of incorrect explanations for scenarios when the AI advice is correct and identifies the negative repercussions for humans.
@@
Taking a human-centric perspective. In our study, we could see that the human-AI team performance improved over the participants’ initial task performance. This pattern, also observed by prior research (Inkpen et al., 2023; Hemmer et al., 2024), showcases the potential of humans collaborating with AI. Even though the scenario in which the human-AI team performance exceeds the performance of human or AI alone—also referred to as complementary team performance (Bansal et al., 2019)—could not be reached (due to the study design choices this was not feasible as the AI advice was always correct), our findings showcase the two sides of explanations: while correct explanations improve the human-AI team performance compared to only receiving AI advice (bright side) and demonstrate the merits of XAI, incorrect explanations decrease the human-AI team performance compared to only receiving AI advice (dark side) and outline the pitfalls of XAI. Thus, it is important to implement mechanisms that allow humans to verify the correctness of explanations, for instance, through reflection mechanisms (Ehsan and Riedl, 2024). Furthermore, we could also identify AI trust as a influencing factor during the main task and cognitive load during the post-test. While these findings align with prior research (Ueno et al., 2022; Buçinca et al., 2021; Westphal et al., 2023), they also emphasize the important role of individuals’ characteristics and traits in human-AI collaboration. Therefore, research has to anticipate these factors in the design of robust and safe explanations.
@@
Anchoring on explanations. We also take a critical view on the phenomenon prior research has identified in collaboration scenarios between humans and AI in which the AI provides explanations: the anchoring effect on explanations (Wang et al., 2019; Bansal et al., 2021). The effect occurs when human decision-makers fixate on the explanation and form an incorrect understanding. In our study, participants in the incorrect explanation treatment demonstrated this behavior by achieving lower procedural knowledge on the post-test compared to the other AI-supported treatments and also by showing lower reasoning capabilities than participants in the control group. As a result, they made the right decisions for the wrong reasons, a condition often referred to as the Clever Hans Effect. (Schramowski et al., 2020; Kauffmann et al., 2020). Although research has adopted the term to describe AI behavior, in our study, humans demonstrate this effect. While this illustrates a major incision in the human knowledge structure, it is crucial for research and practice to develop mechanisms to counteract this effect.
@@
Design guidelines for practice. Our findings confirm that AI support enhances immediate task performance, even when the explanations are incorrect. Even though this observation aligns with prior HCI studies demonstrating that explanations can increase human-AI team performance, such as those by Bansal et al. (2019) and Schemmer et al. (2023b), we extend prior research by revealing the negative impacts of explanations. Our work shows that the performance boost provided by AI during the collaboration can be short-term if the explanations are incorrect, as they negatively impact humans’ procedural knowledge and reasoning post-collaboration. Based on our analyses, we can derive further design implications. Developers of XAI tools should consider how to provide explanations to decision-makers and control for a limited amount of cognitive load infused by the information provided to its users. Our findings also indicate that explanations are necessary to increase trust. However, with the correctness of explanations having impacts on participants procedural knowledge and reasoning, explanations need to be developed and integrated with caution. Additionally, humans with a higher perceived AI usefulness seem to better reason, even for incorrect explanations. This suggests that the design of explanations should be tailored to individuals’ characteristics (e.g., incentivized approaches) in terms of AI usefulness. If participants perceive the AI as useful, they establish higher reasoning abilities. This finding advances the existing literature on XAI, which has primarily emphasized factors like transparency, reliance, and user understanding during the collaboration(Lai and Tan, 2019; Ehsan et al., 2018; Morrison et al., 2024). Our research indicates that while correct explanations foster procedural knowledge, incorrect explanations can lead to the erosion of knowledge over time, even when they temporarily enhance task performance in collaboration with AI. Yet, we note that these implications have to be taken with caution. Even though we illustrate that incorrect explanations have the potential to mislead humans, we cannot generalize for every decision-making scenario. Depending on the context and domain, the impact of incorrect explanations might vary, and organizations have to implement prevention mechanisms designed for the specific scenarios. The findings also underscore the importance of designing AI that prioritize the correctness of explanations, especially in contexts where AI support is not always provided and human decision-makers have to be able to decide autonomously, for instance, as required by the EU AI act (Commission, 2021). We suggest that AI should integrate mechanisms allowing humans to evaluate the correctness of explanations critically. For instance, incorporating counterfactual explanations, as explored by previous research (Wachter et al., 2017; Goyal et al., 2019; Schemmer et al., 2023a), could help users better understand the AI’s decision-making process and mitigate the adverse effects of incorrect explanations. Similarly, maintaining the human ability to perform tasks autonomously without AI support is favorable at the workplace. The role of XAI can impact the transfer of knowledge to counter risks of demographic change (Spitzer et al., 2024b). Incorrect explanations impairing knowledge development have the potential to endanger the successful business operations of organizations. Therefore, we advise designers to incorporate not only mechanisms to detect incorrect explanations post-hoc (Xu et al., 2023; Zhou et al., 2023; Hartwig et al., 2023) but also in the design of the AI.
@@
The impact for organizations. Prior research has mainly explored effects that occur during the collaboration of humans and AI. For instance, Schemmer et al. (2023b) conceptualize the relationship between the appropriateness of reliance and how it relates to the human-AI team performance. Morrison et al. (2024) advance this view by the dimension of explanations and explore the effects of their correctness on humans’ appropriate AI reliance. By outlining potential downsides of XAI, this work addresses the impact of explanations’ correctness on humans post-collaboration. Our study demonstrates that their procedural knowledge and reasoning are impaired when they are provided with incorrect explanations and the AI support is removed. Taking in a human-centric perspective (Horvatić and Lipic, 2021), this repercussion presents harm to not only humans’ individual knowledge development (Bhatt, 2000) but also to their ability to provide meaningful assets to organizations (Nonaka and Takeuchi, 2007; Davenport, 1998). Maintaining individuals as valuable assets to organizations is crucial because it directly influences organizational innovation, efficiency, and adaptability in a competitive market. Schemmer et al. (2021) emphasize that the design of decision-support systems can encourage automation bias and, consequently, deskill human users. Sustaining humans’ knowledge development can foster a more resilient and informed workforce capable of driving sustained success (Grant, 1996).
@@ 6.3. Limitations and Future Work
Despite the valuable insights gained from this study, several limitations must be acknowledged, offering avenues for future research. First, the study explores how incorrect explanations affect humans in AI-assisted decision-making. To investigate the impact of such explanations, we designed a study with four treatments of different AI support: no support, only AI advice, AI advice with correct explanations, and AI advice with incorrect explanations. In real-world applications, the interaction with AI that provides only correct or incorrect explanations is rather unlikely. It presents valuable means to take the first steps to investigate the impact of incorrect explanations but does not reflect the real world. Future research could take on this aspect to extend our findings and evaluate how a mix of correct and incorrect explanations—a mix that is more realistic for deployed AI—affects humans’ procedural knowledge and reasoning. In particular, different ratios of the correctness of explanations could provide further insights and advance the field.
On top of that, our results might not generalize for every human-AI collaboration scenario. We set up a study and gain empirical insights for a decision-making task in which we deploy an LLM to support humans in classifying architectural styles. Even though this task resembles various real-world use-cases in which generative AI is employed to provide advice and additional reasoning (e.g., patients seeking advice for making an initial medical diagnosis or receiving advice for financial investments), we cannot generalize our findings for every context. In different modalities with other task objectives (e.g., content creation) and conditions (e.g., duration of the study), incorrect explanations might show slightly different effects.
@@
Second, the focus of the study was on measuring task performance to derive insights into humans’ procedural knowledge and inform about the impact on human-AI team performance. In addition, we used an LLM-based approach to derive the reasoning abilities of participants. First, such an approach to evaluate the findings poses certain challenges, as LLMs might incorporate biases (Chen et al., 2024; Shi et al., 2024). Second, in real-world scenarios, performance might not be the only metric relevant. Other measures, like appropriate reliance (Schemmer et al., 2023b) or fairness (Schoeffer et al., 2024) in the AI, might also be of high relevance in AI-assisted decision-making as previous studies show (Hemmer et al., 2023; Bansal et al., 2021). It is of high importance to explore how these factors change over time and under the effect of incorrect explanations. Exploring the temporal implications can extend the views and offer new insights that support the robust and effective design of AI.
@@
Lastly, the study primarily relied on short-term measures based on task performance, which may not fully capture the long-term impact of AI support on human knowledge and how the correctness of explanations impacts the human-AI team performance. In our work, participants take, on average, 15 minutes to conduct the study. While such an interaction with AI reflects common real-world scenarios where users briefly consult AI systems for immediate decision support (e.g., seeking advice for math equations in homework or requesting a classification for quality checks), it limits our ability to observe longer-term effects. The insights we gained even in this brief interaction suggest that incorrect explanations can rapidly influence human understanding, though future work should examine whether these effects persist or evolve over extended periods of engagement. The short-term nature of our study may actually underestimate the full impact of incorrect explanations, as longer exposure could lead to deeper entrenchment of misunderstandings. Future research could employ longitudinal designs to assess how incorrect explanations influence procedural knowledge development and reasoning over extended periods and across multiple tasks. This approach would offer a deeper understanding of how different types of explanations contribute to sustained knowledge development, aligning with the principles of human-centered AI.